Solve the integral of √(49-x²)÷x from ln1 to ln2


1
Expert's answer
2021-06-15T04:25:30-0400

"\int_{ln1}^{ln2} {\frac{\sqrt{49-x²}}{x}}dx"

Let x = 7sin"\theta"

dx = 7cos"\theta" d"\theta"

So "\int {\frac{\sqrt{49-x²}}{x}}dx" =

"\int {\frac{\sqrt{49-49sin²\theta}}{7sin\theta}}7cos{\theta}d\theta"

= "\int {\frac{7cos²{\theta}}{sin\theta}}d\theta"

= "7\int {\frac{1-sin²{\theta}}{sin\theta}}d\theta"

= "7\int [{cosec\theta+ sin\theta}]d\theta"

Let ln(cosec"\theta" + cot"\theta") = t

So "\frac{-cosec\theta cot\theta-cosec²\theta)}{cosec\thetaθ + cot\thetaθ)}d\theta"= dt

=> cosec"\theta" d"\theta = -dt"

= "-7\int dt -7\int sin\theta d\theta"

= -7t + 7cos"\theta" + C

= -7 ln|cosec"\theta" + cot"\theta" | + 7cos"\theta"

= 7 ln| "\frac{sin\theta}{1+cos\theta}" | + 7cos"\theta" + C

= "ln[\frac{2sin\theta/2cos\theta/2}{2cos²\theta/2}]+ 7cos\theta + C"

= ln[tan "\theta/2]+ 7cos\theta+C"

Now as x--> ln1 , "\theta" --> 0 and as x--> ln2 , "\theta" --> sin -1(ln2/7)

So

"\int_{ln1}^{ln2} {\frac{\sqrt{49-x²}}{x}}dx" = "[ln(tan\theta/2)+7cos\theta]]_{0}^{sin^{-1}(ln2/7)}"


= "[ln(tan(\frac{1}{2}sin^{-1}(ln(2)/7)+"

"7cos((sin^{-1}(ln(2\/7)]-"

"[ln(tan(0)+" "7cos(0)]"

= "[ln(tan(\frac{1}{2}sin^{-1}(ln(2)/7)+"

"7cos((sin^{-1}(ln(2\/7)] -7 - ln(0)"

= A finite quantity + ∞ as ln(0) -->-∞

= ∞